a machine was used to lift a body of mass of 50 metric tones through a height of 185 meters in 15 minutes. calculate the power of the crane
1
Expert's answer
2018-06-26T13:44:08-0400
Power = Work / Time
Work = weight x height
weight = mass x (gravitational field strength)

Power = mgh / t = 50,000 * 9.81 * 185 / (15 * 60) = 378,094 W or about 378.1 kW